4. (15 pts) Hip dysplasia is a hip socket abnormality that affects many large breed dogs. A review of medical records of dogsPlease show me the R outputs with it
Answer #1

(a)

From table:

P(Hip dysplasia in golden retriever) = 3995/ 46941

P(Hip dysplasia in border collie) = 221/ 5228

The relative risk is:

RR =P(Hip dysplasia in golden retriever) / P(Hip dysplasia in border collie) = ( 3995/ 46941) / ( 221/ 5228) = 2.01

Since RR is greater than 1 so it shows that risk of hip dysplasia is higher in golden retriever group.

(b)

Odds for golden retriever group:

0=3995:42946

Odds for border collie group:

02 = 221 : 5007

Odds ratio:

OR 3999/42946 221/5007 = 2.11

(c)

The 95% confidence interval for odds ratio

In ort106. / 1 In(OR)+1.96/ V 3995 1 1 1 221 +42946 + 5007 = ln (2.11)+1.96 V 3995 +221 + 429465007

= 0.7467 +0.1386 = 0.6081.0.8853)

So required 95% confidence interval for odds ratio is

(0.0081 0.8853) = (1.84, 2.42)

(d)

Since confidence interval does not contain 1 so we cannot conclude that risk of hip dysplasia is same in both group.

----------------------------------------

Following is the R script:

#Given information: a = 3995 b = 42946 C = 221 d = 5007 #part a: Risk in group golden retriever RR1 = a /(a+b) # Risk in grou

Following is the output:

> #Given information: > a = 3995 > b = 42946 > c = 221 > d = 5007 V V V > #part a: Risk in group golden retriever > RR1 = a /